# A Comunication Framework for Compositional Generation

### v1

@@ C.2 Reconstruction games in generative tasks
C.3 The problem reduces to the case of a deterministic student
@@
C.4 Straight-Trough Gumbel-Softmax (STGS) trick
As in (Havrylov & Titov, 2017) and subsequent works, the optimization of policy $\pi_{\varphi}$ (minimization over $\varphi$) can in theory be done via policy gradient optimization methods such as REINFORCE (Williams, 1992), however, due to the combinatorial explosion of the number of messages this can be very compute-intensive if calculated directly, or produces high variance estimators. We thus follow the straight-trough Gumbel-softmax trick (Jang et al., 2017), in which the probabilistic interpretation of $\pi_{\varphi}$ is implicit in the network architecture, and the architecture itself is deterministic and directly differentiable. In this case, we can thus pass to the deterministic description as $m=S_{\phi}(x)$, replacing the probabilistic one $m\sim\pi_{\varphi}(\cdot|m)$, and we obtain the updated form of (37) as
@@ C.4 Straight-Trough Gumbel-Softmax (STGS) trick
C.5 Further ad-hoc improvements to the loss
The actual further improvements of (2) compared to (38) are twofold, and for them our motivation is to bias the model towards higher compositional structure of our message encoding:
  1. 1.
    We use an LSTM architecture for $S_{\phi}$, in which the later words of message $m$ have positional dependency on earlier words. Furthermore, as information is fed back to the model in successive LSTM steps, by the signal processing mutual information inequality, as observed within the Information Bottleneck Principle (Tishby et al., 2000; Shwartz-Ziv & Tishby, 2017) earlier words of a message will have higher mutual information with the input than later words. This signal processing mutual information inequality is a first way in which we bias the model towards condensing higher information content in the first words of the messages.
  2. 2.
    As the space of possible messages is huge, and keeping in mind the simplicity bias of neural networks (Valle-Perez et al., ), we include the geometric series penalizations $\lambda^{i}$ with $10<\lambda<$ 1 multiplying the reconstruction loss of length-$i$ initial subword of $m=S_{\phi}(x)$, and we gradually increase the value of $\lambda$ during training epochs. This process, a simple version of simulated annealing (Kirkpatrick et al., 1983), favors the model fitting the initial words of the messages at the beginning of training, and then gradually improving over the later words of the messages, as the coefficients $\lambda^{i}$ multiplying their contributions to the loss get higher.
Appendix D Pooling Examples
In this section we present two illustrative examples related to the statistics of pooling equilibria. We restrict to the simpler case of $\mathcal{G}=\mathcal{X}=\mathcal{D}$ and $\mathsf{GenX}=id$.
We say that an encoding is locally unambiguous (synonymous to being a pooling solution to the LSG) if it is decodable (i.e. injective) over the training set $\mathcal{D}_{train}$, and globally unambiguous if it is unambiguous over the whole of $\mathcal{D}$ (satisfies UE).
The two considered examples involve training sets $|\mathcal{D}_{train}|$ that differ by a factor of $2$, but they show that the proportion of numbers of global equilibria amongst local equilibria (i.e. the proportion of solutions to the UE condition over $\mathcal{D}_{train}$ that are also UE for $\mathcal{D}$) changes much more drastically, by a factor of $>10$ in this case.
Given a data distribution parameterized by three underlying factors with two values each, specifically let:
[…]
Example 1: easier compositional task. We will take a subset of the distribution as training data, and leave 2 configurations for evaluation, similar to an extrapolation problem (Schott et al., 2022). We will allow a channel capacity of 3 (i.e. take $C=3$) and a binary vocabulary:
[…]
Even though not all combinations are present in the training set, all generating factors are present. This situation becomes very likely when $|\mathcal{D}|$ becomes large, as proved in §A (see the simpler Prop. A.2 or equivalently Thm. A.3 with $k=1$).
We will define a encoding $S$ as a mapping from $\mathcal{X}$ (here $=\mathcal{D}$) to $[V]^{C}$ and a decoding $R$ as a mapping from $[V]^{C}$ to $\mathcal{D}$.
If we do not take into account decodability, encoding amounts to assigning some string in $\mathcal{M}=[V]^{C}$ to each state in $\mathcal{X}=\mathcal{D}$, there are $|\mathcal{D}|^{|V|^{C}}=8^{8}=16777216$ posible encodings.
For an encoding to be decodable without loss, $S$-encoding must form unique pairs from $\mathcal{D}$ to $[V]^{C}$ (condition UE in our setting). For local unambiguity it suffices that all training examples form unique pairs. Then, the amount of codings that are locally unambiguous is:
[…]
The amount of unique pairs from $\mathcal{D}_{train}$ to $[V]^{C}$ is the amount of injective mappings from $\mathcal{D}_{train}$ to $[V]^{C}$, i.e, $\frac{[V]^{C}!}{([V]^{C}-|\mathcal{D}_{train}|)!}=20160$. The amount of encodings from $\mathcal{D}\setminus\mathcal{D}_{train}$ to $[V]^{C}$ is $(|\mathcal{D}|-|\mathcal{D}_{train}|)^{|V|^{C}}=2^{8}=256$. Therefore the amount of locally optimal encodings is $20160*256=5160960$. The number of globally optimal encodings is $\frac{[V]^{C}!}{([V]^{C}-|\mathcal{D}|)!}=40320$. This implies that for this problem, the ratio of globally unambiguous to locally unambiguous solutions is $7,8125\times 10^{-3}$.
Example 2: hard compositional task. For the same $\mathcal{D}$ as above, we take
$\mathcal{D}_{train}:=\{\{\square,\rightarrow,\uparrow\},\{\circ,\leftarrow,\downarrow\}\},$
maintaining $C=3$ and $V=\{0,1\}$ as in the previous example.
This is, in a sense, ”the hardest possible” amongst solvable compositional reconstruction tasks, in the sense that this choice of $\mathcal{D}_{train}$ has minimum cardinality amongst training set selections such that all factor values are exemplified in some of their elements.
Compared to the previous example, the set of possible encodings and the set of globally unambiguous encodings remain the same, however, the amount of locally unambiguous mappings increases drastically compared to the previous example, and now is $\frac{[V]^{C}!}{([V]^{C}-|\mathcal{D}_{train}|)!}*|\mathcal{D}\setminus\mathcal{D}_{train}|^{|V|^{C}}=94058496$, and thus the ratio of globally to locally unambiguous encodings decreases drastically to $4,28669\times 10^{-4}$
This illustrates the phenomenon that the ratio of global optima vs pooling solutions is very sensitive to the richness of $\mathcal{D}_{train}$.
Appendix E Hyperparameter search
Beta Lambda $\ell_{10^{-3}}$ Top Sim $\uparrow$ Final Symbol Loss $\downarrow$ Test Top Sim $\uparrow$ Test Final Symbol Loss $\downarrow$
0.5 0.75 4.7076 0.1802 0.0277 0.2333 0.0425
1 0.75 4.8052 0.1547 0.0512 0.2152 0.0520
10 0.75 4.9126 0.1597 0.0440 0.2156 0.0534
  0.5 0.95 5.2227 0.1894 0.0306 0.2681 0.0391
  1 0.95 5.3032 0.1761 0.0359 0.2335 0.0457
10 0.95 5.5674 0.1829 0.0436 0.2382 0.0456
0.5 1 5.3263 0.1766 0.0348 0.2526 0.0385
1 1 5.4198 0.1706 0.0362 0.2357 0.0492
10 1 5.5297 0.1661 0.0312 0.2284 0.0534
Table 2: Comparison of Hyperparameter Combinations (averaged over 10 runs), with first three metrics calculated at epoch 200 on training set, and the last two calculated at test after 200-epoch training. Recall that $\ell_{\epsilon}$ is the minimum message length, averaged over the dataset, here used with thresholding value $\epsilon=10^{-3}$. While low values of $\ell_{\epsilon}$ are better towards good compression (EE), this is only true assuming good accuracy, and otherwise small $\ell_{\epsilon}$ values may be unfavorable. TopSim abbreviates topological similarity. We highlight the combination used in the graphs in the body of the paper, $\beta=0.5$, $\lambda=0.95$, which, although not optimal for all metrics, is optimal for TopSim, and a good compromise on the other metrics.
Appendix F Generation examples
F.1 Reconstruction
@@
The following reconstruction examples were generated with the ILFI+PDM model trained for 200 epochs, with $\lambda=0.95$ and $\beta=0.5$. From left to right and top to bottom, the first two images of each example are the original image to be reconstructed, and the reconstruction generated by the VAE backbone. The following images are the reconstructions produced by evaluated the predicted vector $z_{i}=R_{\omega}(S_{\phi}(x)_{[i]})$ with the VAE backbone.
@@
We observe semantic differences between each prediction (flor hue, shape, object color, angle, etc.), as opposed to increasing in image quality or clarity. This fact leads us to believe that the message space effectively captures the generating factors $\mathcal{G}$ and is likely to be compositional in some measure.
Figure 4: Reconstructed images.
F.2 Interpolation
A possible way to evaluate the compositionality within the learned message space is to interpolate between representations of different objects in this space. Specifically, given two states $x_{1}$ and $x_{2}$, we visualize iteratively converting the message from one into the other.
In the following images, the first row consists of the original states $x_{1}$ (left) and $x_{2}$ (right). From the second row onwards, If $m^{1}=S_{\phi}(x_{1})$ and $m^{2}=S_{\phi}(x_{2})$, the $i_{th}$ row is the reconstruction from the message formed by concatenating $m^{1}_{[1:C-i]}$ and $m^{2}_{[C-i:C]}$, starting from 0 (i.e the second row from the top is simply $R_{\omega}(m^{1})$). We find that, as opposed to smoothly interpolating in generative model latent spaces, this exercise produces interpretable discrete steps to semantically interpolate the states. We invite the reader to observe the diagonal (starting from the rightmost cell on the third row), as this is where the message changes, and make note that the position of the diagonal coincides with some semantic change in the reconstruction.
Figure 5: Semantic interpolation 1.
Figure 6: Semantic interpolation 2.
Figure 7: Semantic interpolation 3.
Figure 8: Semantic interpolation 4.
Figure 9: Semantic interpolation 5.
Figure 10: Semantic interpolation 6.

### v3

@@ G.2 Reconstruction games in generative tasks
G.3 The problem reduces to the case of a deterministic student
@@
G.4 Straight-Trough Gumbel-Softmax (STGS) trick
As in [18] and subsequent works, the optimization of policy $\pi_{\varphi}$ (minimization over $\varphi$) can in theory be done via policy gradient optimization methods such as REINFORCE [58], however, due to the combinatorial explosion of the number of messages this can be very compute-intensive if calculated directly, or produces high variance estimators. We thus follow the straight-trough Gumbel-softmax trick [19], in which the probabilistic interpretation of $\pi_{\varphi}$ is implicit in the network architecture, and the architecture itself is deterministic and directly differentiable. In this case, we can thus pass to the deterministic description as $m=S_{\phi}(x)$, replacing the probabilistic one $m\sim\pi_{\varphi}(\cdot|m)$, and we obtain the updated form of (47) as
@@ G.4 Straight-Trough Gumbel-Softmax (STGS) trick
Appendix H Hyperparameter search
Table 3: Comparison of TopSim, Final MSE, and Useful Length for different regularizations and values of $\beta$ and $\lambda$.
Dataset Mode Lambda Beta TopSim Final MSE Useful Length
Shapes3D PDM 1.0 1.25 0.264 $\pm$ 0.010 0.208 $\pm$ 0.020 7.600 $\pm$ 0.780
PDM 1.0 1.5 0.295 $\pm$ 0.010 0.182 $\pm$ 0.020 6.600 $\pm$ 0.610
PDM 5.0 1.25 0.262 $\pm$ 0.020 0.215 $\pm$ 0.030 7.600 $\pm$ 0.780
PDM 5.0 1.5 0.260 $\pm$ 0.010 0.178 $\pm$ 0.010 5.000 $\pm$ 0.280
PDM 10.0 1.25 0.279 $\pm$ 0.010 0.191 $\pm$ 0.020 6.200 $\pm$ 0.660
PDM 10.0 1.5 0.276 $\pm$ 0.010 0.177 $\pm$ 0.010 6.400 $\pm$ 0.610
KoLeo 1.0 1.25 0.222 $\pm$ 0.020 0.272 $\pm$ 0.040 8.200 $\pm$ 0.720
KoLeo 1.0 1.5 0.265 $\pm$ 0.000 0.184 $\pm$ 0.010 7.000 $\pm$ 0.630
KoLeo 5.0 1.25 0.258 $\pm$ 0.010 0.198 $\pm$ 0.000 7.800 $\pm$ 0.660
KoLeo 5.0 1.5 0.247 $\pm$ 0.020 0.222 $\pm$ 0.020 8.000 $\pm$ 0.570
KoLeo 10.0 1.25 0.244 $\pm$ 0.010 0.232 $\pm$ 0.030 7.500 $\pm$ 0.550
KoLeo 10.0 1.5 0.257 $\pm$ 0.010 0.202 $\pm$ 0.010 6.800 $\pm$ 0.500
MPI3D PDM 1.0 1.25 0.154 $\pm$ 0.000 0.020 $\pm$ 0.000 6.200 $\pm$ 1.530
PDM 1.0 1.5 0.160 $\pm$ 0.000 0.020 $\pm$ 0.000 6.600 $\pm$ 1.310
PDM 5.0 1.25 0.152 $\pm$ 0.000 0.020 $\pm$ 0.000 5.000 $\pm$ 1.470
PDM 5.0 1.5 0.158 $\pm$ 0.010 0.020 $\pm$ 0.000 6.800 $\pm$ 1.210
PDM 10.0 1.25 0.161 $\pm$ 0.000 0.020 $\pm$ 0.000 6.200 $\pm$ 1.530
PDM 10.0 1.5 0.162 $\pm$ 0.000 0.020 $\pm$ 0.000 7.800 $\pm$ 1.070
KoLeo 1.0 1.25 0.162 $\pm$ 0.000 0.020 $\pm$ 0.000 4.000 $\pm$ 1.170
KoLeo 1.0 1.5 0.160 $\pm$ 0.000 0.020 $\pm$ 0.000 7.600 $\pm$ 1.250
KoLeo 5.0 1.25 0.155 $\pm$ 0.000 0.020 $\pm$ 0.000 6.600 $\pm$ 1.340
KoLeo 5.0 1.5 0.163 $\pm$ 0.000 0.020 $\pm$ 0.000 5.800 $\pm$ 1.210
KoLeo 10.0 1.25 0.160 $\pm$ 0.000 0.020 $\pm$ 0.000 4.000 $\pm$ 1.170
KoLeo 10.0 1.5 0.159 $\pm$ 0.000 0.020 $\pm$ 0.000 5.800 $\pm$ 1.180
Appendix I Generation examples
I.1 Reconstruction
@@
The following reconstruction examples were generated with the CELEBI model, with PDM regularization, $\beta=1$ and $\lambda=1.5$. From left to right, the first image of each example is the original image to be reconstructed, and the following images are the predicted reconstructions $x_{i}=R_{\omega}(S_{\phi}(x)_{[1:i]})$ with the VAE backbone.
@@
We observe semantic differences between each prediction (floor hue, shape, object color, angle, etc.), as opposed to increasing in image quality or clarity. This fact leads us to believe that the message space effectively captures the generating factors $\mathcal{G}$ and is likely to be compositional in some measure.
(a) Example 1
(b) Example 2
(c) Example 3
(d) Example 4
Figure 5: Progressive image reconstructions obtained at each decoding step by the receiver as it processes successive symbols of the message.